# Weak measurement-based protocol for ergotropy protection in open quantum batteries

### v1

@@ III.1 Incoherent ergotropy
Figure 2: Incoherent ergotropy gain as a function of the weak measurement strength $m$ and initial population $P_{0}$ along the null energetic shift reversal measurement strength $\tilde{w}\in[0,1]$ for distinct dissipation times (a) $\tau=0.5\tau_{\gamma}$, (b) $\tau=\tau_{\gamma}$, and (c) $\tau=10\tau_{\gamma}$. The red-hatched region and red curves indicate that $\mathcal{W}_{m\tilde{w}}(\tau)=0$. The $\eta_{2}$ lines are highlighted.
@@
Along with null energy and ergotropy shifts, the protocol performance is characterized by the ergotropy gain, quantifying the amount of ergotropy saved, relative to a quantum battery subjected to a purely discharging process. In this context, Fig. 2 illustrates the behavior of incoherent ergotropy gain as a function of the weak measurement strength $m$ and the initial population $P_{0}$, computed along $\tilde{w}\in[0,1]$ at various dissipation times: (a) $0.5\tau_{\gamma}$, (b) $\tau_{\gamma}$, and (c) $10\tau_{\gamma}$. For clarity, the region where ¯$\overline{\mathcal{R}}_{m\tilde{w}}^{inc}(\tau)\leq 0$—indicating no gain following protocol execution—has been omitted. The red-hatched areas and red curves denote operational regions where $\mathcal{W}_{m\tilde{w}}(\tau)=0$. The red lines $\eta_{2}$ show points characterized by the equal-strength curve, as defined in Eq. (25). Notice that, in general, stronger measurements $m$ yield higher gains. It is also evident that not all parameter space corresponds to valid operational regions. The size of these working regions depends strongly on the dissipation time $\tau$, gradually decreasing over time and eventually saturating for long dissipation times, characterized by the reversal measurement strength $\lim_{\tau\rightarrow\infty}\tilde{w}$. In this regime, the only remaining valid execution of the protocol with a finite gain is characterized by $\eta_{2}$.
@@
To better illustrate how the protocol functions, Fig. 3 shows the variation of the incoherent ergotropy during each step (dashed-blue line), assuming $P_{0}=0.9$, $m=0.4$, dissipation time $\tau=\tau_{\gamma}$, and $w=\tilde{w}\approx 0.2$, which ensure $\varepsilon_{m\tilde{w}}(\tau_{\gamma})=0$. This is compared to the purely discharging process (continuous-red line). The blue arrows indicate the discontinuous (instantaneous) ergotropy jumps caused by the weak and reversal measurements. According to Fig. 2b, the chosen parameters lie within the operational region, with $\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$. Note that while $\hat{M}_{m}$ decreases the initial battery ergotropy, the execution of the final step, $\hat{W}_{\tilde{w}}$, increases it. Furthermore, the battery becomes nearly fully discharged due to environmental interaction at step (iii), since $\tau_{\gamma}\approx\tau_{1/2}$. As a result, at the end of the protocol, the final incoherent ergotropy $\mathcal{R}^{inc}_{iv}$ exceeds the red curve, yielding ¯$\overline{\mathcal{R}}_{m\tilde{w}}^{inc}(\tau_{\gamma})\approx 0.71\times 10^{-1}$. It is important to emphasize that this gain is achieved without actively introducing any additional external resources, which is ensured by an appropriate choice of parameters that results in null energetic and ergotropic shifts.
@@ III.2 Coherent ergotropy
Figure 4: Coherent ergotropy gain as a function of $m$ and initial coherence $Q_{0}\in[0,\sqrt{P_{0}(1-P_{0})}]$ computed along $\tilde{w}\in[0,1]$ and $\tau_{\gamma}$ for (a) $P_{0}=0.4$, (b) $P_{0}=0.6$, (c) $P_{0}=0.9$. The dashed blue line represents the points where ¯$\overline{\mathcal{R}}^{coh}_{m\tilde{w}}(\tau_{\gamma})=0$. The red lines highlight the region where $\mathcal{W}_{m\tilde{w}}(\tau)=0$.
@@
Figure 5: Coherent ergotropy gain in terms of dissipation time and measurement strength $m$ for (a) $P_{0}=0.6$ and (b) $P_{0}=0.9$, assuming $Q_{0}=Q_{max}$ and $\tilde{w}\in[0,1]$. The blue dashed lines represent the points with null coherent ergotropy gain. The red curves indicate the operational points, where $\mathcal{W}_{m\tilde{w}}(\tau)=0$.
@@
The role played by the coherence for the protocol is fully encoded into Eq. (15). Since the measurement operations are incoherent, this quantity is non-null iff $Q_{0}\neq 0$. Hence, given $\hat{\rho}_{0}$, the initial coherent component of the ergotropy is written as
$\mathcal{R}_{i}^{coh}=\frac{1}{2}\left(-|1-2P_{0}|+\sqrt{(1-2P_{0})^{2}+4|Q_{0}|^{2}}\right).$ (32)
@@
After performing the weak measurement at step (ii), the ergotropy jumps to
$\begin{split}\mathcal{R}_{ii}^{coh}&=-\frac{|2-2P_{0}-\textrm{N}_{m}|}{2\textrm{N}_{m}}\\ &+\frac{\sqrt{(2-2P_{0}-\textrm{N}_{m})^{2}-4|Q_{0}|^{2}\frac{(1-P_{0}-\textrm{N}_{m})}{P_{0}}}}{2\textrm{N}_{m}}.\end{split}$ (33)
In contrast to the incoherent counterpart, the behavior of coherent ergotropy during the interaction with the environment is more complex. See Appendix A for more details. Specifically, the coherent ergotropy does not necessarily decay monotonically over time, even though $Q_{m}(t)$ exhibits monotonic behavior. Therefore, during step (iii), one has
$\begin{split}\mathcal{R}_{iii}^{coh}&=-\frac{1}{2}|1-2P_{m}(t)|\\ &+\frac{1}{2}\sqrt{(1-2P_{m}(t))^{2}+4|Q_{m}(t)|^{2}}.\end{split}$ (34)
Finally, after the reversal measurement at step (iv), the coherent ergotropy becomes
@@
$\begin{split}\mathcal{R}_{iv}^{coh}&=-\frac{|\textrm{N}_{mw}(\tau)-2P_{m}(\tau)|}{2\textrm{N}_{mw}(\tau)}\\ &+\frac{\sqrt{(\textrm{N}_{mw}(\tau)-2P_{m}(\tau))^{2}+\frac{4|Q_{m}(\tau)|^{2}(P_{m}(\tau)-\textrm{N}_{mw}(\tau))}{\left(P_{m}(\tau)-1\right)}}}{2\textrm{N}_{mw}(\tau)}.\end{split}$ (35)
@@
To provide a broader characterization of the role of coherence, Fig. 4 illustrates the landscape of coherent ergotropy gain in terms of $m$ and initial coherence $Q_{0}\in[0,\sqrt{P_{0}(1-P_{0})}]$. This is computed along $\tilde{w}\in[0,1]$ at $\tau_{\gamma}$ for fixed initial populations: (a) $P_{0}=0.4$, (b) $P_{0}=0.6$, and (c) $P_{0}=0.9$. As before, the red lines indicate the working points, where $\mathcal{W}_{m\tilde{w}}(\tau)=0$. It is evident that, depending on the parameters, coherence can offer an advantage to the protocol and contribute positively to the ergotropy gain. Nevertheless, at $P_{0}=0.6$, no combination of $Q_{0}$ and $m$ yields a ¯$\overline{\mathcal{R}}^{coh}_{m\tilde{w}}(\tau_{\gamma})>0$ at operational points. Still, it is noteworthy that its scale is one order of magnitude lower than the incoherent part, suggesting that the overall protocol may still be beneficial despite small negative contributions. Also, notice that, given the monotonic decay of the coherence independently of the protocol, no gain is anticipated for long dissipation times. In this sense, Fig. 5 complements Fig. 4b and Fig. 4c by showing how the coherent ergotropy gain behaves in terms of $\tau$ for different measurement strengths $m$ and $\tilde{w}\in[0,1]$, assuming maximum real value allowed for each $P_{0}$, given by $Q_{0}=Q_{max}\coloneqq\sqrt{P_{0}(1-P_{0})}$ for (a) $P_{0}=0.6$, (b) $P_{0}=0.9$. It is possible to see that, for small $\tau$, the operational values of $m$ change. After that, it asymptotically approaches a stationary value.
@@
Figure 6: Coherent ergotropy comparison. Coherent ergotropy changes with (dashed blue) and without (continuous red) the execution of the protocol for the same parameters used in Fig. 3 and $Q_{0}\approx 0.28$. The blue arrows highlight the coherent ergotropy jumps due to the measurements.
@@
Finally, as an illustrative example, Fig. 6 displays how the coherent ergotropy changes at each step of the protocol, assuming the same set of parameters used in Fig. 3 with the additional initial coherence $Q_{0}\approx 0.28$ to achieve an operational point (see Fig. 4c). As previously mentioned, this quantity exhibits non-trivial behavior during the environmental interaction. Namely, one can show that for $P_{0}>1/2$, $\mathcal{R}_{iii}^{coh}$ increases over time, reaching a turning point at $\tau_{1/2}$ given by Eq. (31), when the diagonal state becomes passive. Given that $\tau_{\gamma}\approx\tau_{1/2}$ for the chosen parameters, such non-monotonicity is not observed. Also, in contrast to the incoherent part, the action of $\hat{M}_{m}$ and $\hat{W}_{\tilde{w}}$ respectively increase and decrease the coherent ergotropy. Despite that, at the end of the protocol $\mathcal{R}_{iv}^{coh}$ still remains above the red curve, resulting in a small finite gain of ¯$\overline{\mathcal{R}}^{coh}_{m\tilde{w}}(\tau_{\gamma})\approx 0.5\times 10^{-2}$.
@@ III.3 Total ergotropy gain and success probability
Figure 7: Protocol performance in terms of the weak measurement strength $m$ and initial population $P_{0}$. These quantities were computed along the null energetic shift curve $\tilde{w}\in[0,1]$, considering $Q_{0}=Q_{max}$ for each $P_{0}$, and $\tau=\tau_{\gamma}$. The dashed blue curves indicate the points with null ergotropy gain, while the red lines highlight the points satisfying $\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$. (a) Total ergotropy gain; (b) Coherence contribution to the protocol, quantified by the coherent ergotropy gain; (c) Protocol’s success probability.
@@
Figure 8: Success probability in terms of the ergotropy gain and $P_{0}$ for $m=0.1$, $m=0.5$ and $m=0.9$, computed at $\tau_{\gamma}$, $\tilde{w}\in[0,1]$ and $Q_{0}=Q_{max}$ for each $P_{0}$. The red dots represent the operational points of each curve, where $\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$.
Figure 9: Ergotropy comparison. Ergotropy changes with (dashed blue) and without (continuous red) the execution of the protocol for $P_{0}=0.9$, $Q_{0}\approx 0.28$, dissipation time $\tau=\tau_{\gamma}$, and strengths $m=0.4$ and $w=\tilde{w}\approx 0.20$. The blue arrows highlight the ergotropy jumps due to the measurements. The probability of success is $\Pi_{m\tilde{w}}(\tau_{\gamma})\approx 0.57$.
@@
To better assess the general performance of the protocol, Fig. 7 presents (a) the total and (b) coherent ergotropy gains landscape in terms of the initial population $P_{0}$ and the weak measurement strength $m$, and (c) the probability of success. These quantities were computed along $\tilde{w}\in[0,1]$, dissipation time $\tau=\tau_{\gamma}$ and for the initial coherence it was assumed $Q_{0}=Q_{max}$ for each $P_{0}$. The blue dashed lines represent the points where the ergotropy gains are null, while the red continuous line highlights the protocol’s working points, where $\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$ is also satisfied. Notice that, despite the total ergotropy gain being positive for most of the $P_{0}\times m$ surface, for the chosen set of parameters the protocol has no operational points above a certain value of $P_{0}\approx 0.85$. Also, it is possible to observe that one obtains higher gains for stronger measurements, even in regions where the coherence contribution is negative. As a probabilistic protocol, its performance should be evaluated alongside its probability of success. In this sense, notice that ¯$\overline{\mathcal{R}}_{m\tilde{w}}(\tau_{\gamma})$ and $\Pi_{m\tilde{w}}(\tau_{\gamma})$ are inversely related, i.e., higher gains are obtained with lower probabilities of success and vice versa. On the other hand, the coherent ergotropy gain has a positive contribution with higher probabilities. It should be stressed that the observed probability landscape is a generic feature of the protocol, due to its independence of the initial coherence $Q_{0}$, i.e., the fact that the probability is maximized in the direction of lower $m$ and $P_{0}$ is observed regardless of the amount of coherence initially assumed. To better understand the relationship between gain and probability, Fig. 8 illustrates the probability of success as a function of the ergotropy gain and $P_{0}$ for measurement strengths $m=0.1$, $m=0.5$, and $m=0.9$, assuming $Q_{0}=Q_{max}$ for each $P_{0}$. The red dots denote the working points of each curve, where $\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$ is also satisfied. Notably, for $m=0.1$, a small gain is achieved with a high probability, whereas for $m=0.9$, the gain is significantly higher but comes at the cost of a much lower probability, indicating an inverse relationship between these two quantities. Therefore, by tuning to intermediate measurement strengths, it is possible to identify operational points that balance finite gain with effective probability.

### v2

@@ III.1 Incoherent ergotropy
Along with null energy and ergotropy shifts, the protocol performance is characterized by the ergotropy gain, quantifying the amount of ergotropy saved, relative to a quantum battery subjected to a purely discharging process. In this context, Fig. 2 illustrates the behavior of incoherent ergotropy gain as a function of the weak measurement strength $m$ and the initial population $P_{0}$, computed along $\tilde{w}\in[0,1]$ at various dissipation times: (a) $0.5\tau_{\gamma}$, (b) $\tau_{\gamma}$, and (c) $10\tau_{\gamma}$. For clarity, the region where ¯$\overline{\mathcal{R}}_{m\tilde{w}}^{inc}(\tau)\leq 0$—indicating no gain following protocol execution—has been omitted. The red-hatched areas and red curves denote operational regions where $\mathcal{W}_{m\tilde{w}}(\tau)=0$. The red lines $\eta_{2}$ show points characterized by the equal-strength curve, as defined in Eq. (25). Notice that, in general, stronger measurements $m$ yield higher gains, being a direct consequence of the smaller discharging rate for less charged states. It is also evident that not all parameter space corresponds to valid operational regions. The size of these working regions depends strongly on the dissipation time $\tau$, gradually decreasing over time and eventually saturating for long dissipation times, characterized by the reversal measurement strength $\lim_{\tau\rightarrow\infty}\tilde{w}$. In this regime, the only remaining valid execution of the protocol with a finite gain is characterized by $\eta_{2}$.
Figure 4: Coherent ergotropy profile in terms of $P$ and $|Q|^{2}\in[0,P(1-P)]$. The black (red) curves represent the dissipative dynamics of Eq. (3) for initial states with $P(0)<1/2$ ($P(0)>1/2$). The purple star highlights the thermal state, while the red dots indicate the instant $\tau_{1/2}$ when $P(\tau_{1/2})=1/2$.
@@
To better illustrate how the protocol functions, Fig. 3 shows the variation of the incoherent ergotropy during each step (dashed-blue line), assuming $P_{0}=0.9$, $m=0.4$, dissipation time $\tau=\tau_{\gamma}$, and $w=\tilde{w}\approx 0.2$, which ensure $\varepsilon_{m\tilde{w}}(\tau_{\gamma})=0$. This is compared to the purely discharging process (continuous-red line). The blue arrows indicate the discontinuous (instantaneous) ergotropy jumps caused by the weak and reversal measurements. According to Fig. 2b, the chosen parameters lie within the operational region, with $\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$. Note that while $\hat{M}_{m}$ decreases the initial battery ergotropy, the execution of the final step, $\hat{W}_{\tilde{w}}$, increases it. Furthermore, the battery becomes nearly fully discharged due to environmental interaction at step (iii), since $\tau_{\gamma}\approx\tau_{1/2}$. As a result, at the end of the protocol, the final incoherent ergotropy $\mathcal{R}^{inc}_{iv}$ exceeds the red curve, yielding ¯$\overline{\mathcal{R}}_{m\tilde{w}}^{inc}(\tau_{\gamma})\approx 0.71\omega\times 10^{-1}$, which represents a saving of $100\times\overline{\mathcal{R}}_{m\tilde{w}}^{inc}(\tau_{\gamma})/\mathcal{R}_{i}\approx 8.89\%$ of the initial charge. It is important to emphasize that this gain is achieved without actively introducing any additional (net) external resources, which is ensured by an appropriate choice of parameters that results in null energetic and ergotropic shifts.
@@ III.2 Coherent ergotropy
Figure 5: Coherent ergotropy gain as a function of $m$ and initial coherence $|Q_{0}|^{2}\in[0,P_{0}(1-P_{0})]$ computed along $\tilde{w}\in[0,1]$ and $\tau_{\gamma}$ for (a) $P_{0}=0.4$, (b) $P_{0}=0.6$, (c) $P_{0}=0.9$. The white line represents the points where ¯$\overline{\mathcal{R}}^{coh}_{m\tilde{w}}(\tau_{\gamma})=0$. The red lines highlight the region where $\mathcal{W}_{m\tilde{w}}(\tau)=0$.
@@
Figure 6: Coherent ergotropy gain in terms of dissipation time and measurement strength $m$ for (a) $P_{0}=0.6$ and (b) $P_{0}=0.9$, assuming $|Q_{0}|^{2}=|Q_{max}|^{2}$ and $\tilde{w}\in[0,1]$. The white dashed lines represent the points with null coherent ergotropy gain. The red curves indicate the operational points, where $\mathcal{W}_{m\tilde{w}}(\tau)=0$.
The effect of quantum coherence in the protocol can be analyzed by employing the coherent ergotropy, defined in Eq. (15). Before analyzing each step, it is useful to highlight some general properties that hold regardless of the protocol. Consider a quantum state of the form $\hat{\rho}=\left(1-P\right)|g\rangle\langle g|+P|e\rangle\langle e|+\left(Q|g\rangle\langle e|+h.c.\right)$. The purity of this state is given by $\mu[\hat{\rho}]=P^{2}+(1-P)^{2}+2|Q|^{2}\leq 1$, which implies that the coherent term is constrained by $\left|Q\right|^{2}\leq P\left(1-P\right)$. From this, the coherent ergotropy is expressed as
$\mathcal{R}^{coh}[\hat{\rho}]=\frac{\omega}{2}\left(-|1-2P|+\sqrt{(1-2P)^{2}+4|Q|^{2}}\right).$ (32)
By construction, $\mathcal{R}^{coh}[\hat{\rho}]$ is always non-negative and vanishes whenever coherence is absent ($Q=0$), including at the limits $P=0$ and $P=1$. Additionally, this function is symmetric with respect to population inversion and coherence conjugation, i.e., $\mathcal{R}^{coh}[\hat{\rho}(P,Q)]=\mathcal{R}^{coh}[\hat{\rho}(1-P,Q)]$ and $\mathcal{R}^{coh}[\hat{\rho}(P,Q)]=\mathcal{R}^{coh}[\hat{\rho}(P,Q^{*})]$, with maximum value $\mathcal{R}^{coh}[\hat{\rho}]=\omega|Q|$ which occurs at $P=1/2$, leading to an increasing behavior in terms of $P$ for $P<1/2$ and a decreasing behavior for $P>1/2$. Furthermore, the convexity condition $\partial^{2}\mathcal{R}^{coh}[\hat{\rho}]/\partial P^{2}>0$ holds for all values of $P$ and $|Q|$.
Figure 4 illustrates the coherent ergotropy as a function of population $P$ and coherence $|Q|^{2}$, highlighting these key properties. Any dynamical evolution follows a trajectory within this landscape. Specifically, under the evolution governed by Eq. (3), the behavior of $\mathcal{R}^{coh}[\hat{\rho}(t)]$ depends on both the initial state $(P(0),|Q(0)|^{2})$ and the stationary thermal state characterized by a fixed population $f\in[0,1/2)$. During thermalization, $|Q(t)|^{2}$ always decays exponentially, while $P(t)$ asymptotically approaches $f$, with $dP(t)/dt>0$ for $0\leq P(0)<f$ and $dP(t)/dt<0$ for $f<P(0)\leq 1$. In Figure 4, black (red) curves represent trajectories for initial states with $P(0)<1/2$ ($P(0)>1/2$). Regardless of the initial condition, the system reaches the stationary state given by $f$ (purple star). When $P(0)<1/2$, $\mathcal{R}^{coh}$ decreases monotonically. However, for $P(0)>1/2$, we observe a more complex behavior: the coherent ergotropy initially increases until the population reaches $P=1/2$ at time $\tau_{1/2}$, after which it begins to decrease (asymptotically reaching $f$). This non-monotonic behavior means that before the inevitable discharge, systems with $P(0)>1/2$ experience a temporary enhancement of their coherent ergotropy during the interval $t\in[0,\tau_{1/2}]$.
@@
Considering the TWM protocol, since the measurement operations are incoherent, the coherent ergotropy is non-null iff $Q_{0}\neq 0$. Hence, given $\hat{\rho}_{0}$ and Eq. (32), the initial coherent component of the ergotropy is written as
$\mathcal{R}_{i}^{coh}=\frac{\omega}{2}\left(-|1-2P_{0}|+\sqrt{(1-2P_{0})^{2}+4|Q_{0}|^{2}}\right).$ (33)
@@
After performing the weak measurement at step (ii), in terms of the relevant parameters, the ergotropy jumps to
$\begin{split}\mathcal{R}_{ii}^{coh}=&-\omega\frac{|2-2P_{0}-\textrm{N}_{m}|}{2\textrm{N}_{m}}\\ &+\omega\frac{\sqrt{(2-2P_{0}-\textrm{N}_{m})^{2}-4|Q_{0}|^{2}\frac{(1-P_{0}-\textrm{N}_{m})}{P_{0}}}}{2\textrm{N}_{m}}.\end{split}$ (34)
After contact with the thermal reservoir, during step (iii), one has
$\begin{split}\mathcal{R}_{iii}^{coh}=&-\frac{\omega}{2}|1-2P_{m}(t)|\\ &+\frac{\omega}{2}\sqrt{(1-2P_{m}(t))^{2}+4|Q_{m}(t)|^{2}}.\end{split}$ (35)
As mentioned above, in contrast to the incoherent counterpart, the transient dynamical behavior of the coherent ergotropy during the thermalization process is considerably less trivial. Thus, depending on the initial state of the dissipation, given by the post-weak measurement state and population $P_{m}(0)$, $\mathcal{R}_{iii}^{coh}$ may exhibit non-monotonic behavior with a temporary increase. Finally, after the reversal measurement at step (iv), the coherent ergotropy becomes
@@
$\begin{split}\mathcal{R}_{iv}^{coh}&=-\omega\frac{|\textrm{N}_{mw}(\tau)-2P_{m}(\tau)|}{2\textrm{N}_{mw}(\tau)}\\ &+\omega\frac{\sqrt{(\textrm{N}_{mw}(\tau)-2P_{m}(\tau))^{2}+4|Q_{m}(\tau)|^{2}\frac{(P_{m}(\tau)-\textrm{N}_{mw}(\tau))}{\left(P_{m}(\tau)-1\right)}}}{2\textrm{N}_{mw}(\tau)}.\end{split}$ (36)
@@
To provide a broader characterization of the role of coherence, Fig. 5 illustrates the landscape of coherent ergotropy gain in terms of $m$ and initial coherence $|Q_{0}|^{2}\in[0,P_{0}(1-P_{0})]$. This is computed along $\tilde{w}\in[0,1]$ at $\tau_{\gamma}$ for fixed initial populations: (a) $P_{0}=0.4$, (b) $P_{0}=0.6$, and (c) $P_{0}=0.9$. As before, the red lines indicate the working points, where $\mathcal{W}_{m\tilde{w}}(\tau)=0$. It is evident that, depending on the parameters, coherence can offer an advantage to the protocol and contribute positively to the ergotropy gain. Nevertheless, at $P_{0}=0.6$, no combination of $Q_{0}$ and $m$ yields a ¯$\overline{\mathcal{R}}^{coh}_{m\tilde{w}}(\tau_{\gamma})>0$ at operational points. Still, it is noteworthy that its scale is one order of magnitude lower than the incoherent part, suggesting that the overall protocol may still be beneficial despite small negative contributions. Also, notice that, given the monotonic decay of the coherence independently of the protocol, no gain is anticipated for long dissipation times. In this sense, Fig. 6 complements Fig. 5b and Fig. 5c by showing how the coherent ergotropy gain behaves in terms of $\tau$ for different measurement strengths $m$ and $\tilde{w}\in[0,1]$, assuming maximum coherence allowed for each $P_{0}$, given by $|Q_{0}|^{2}=|Q_{max}|^{2}\coloneqq P_{0}(1-P_{0})$ for (a) $P_{0}=0.6$, (b) $P_{0}=0.9$. It is possible to see that, for small $\tau$, the operational values of $m$ change. After that, it asymptotically approaches a stationary value.
@@
Figure 7: Coherent ergotropy comparison. Coherent ergotropy changes with (dashed blue) and without (continuous red) the execution of the protocol for the same parameters used in Fig. 3 and $|Q_{0}|^{2}\approx 7.67\times 10^{-2}$. The blue arrows highlight the coherent ergotropy jumps due to the measurements.
@@
Finally, as an illustrative example, Fig. 7 displays how the coherent ergotropy changes at each step of the protocol, assuming the same set of parameters used in Fig. 3 with the additional initial coherence satisfying $|Q_{0}|^{2}\approx 7.67\times 10^{-2}$ to achieve an operational point (see Fig. 5c). As anticipated, since $P_{m}(0)=P_{0}(1-m)/\textrm{N}_{m}>1/2$, the coherent ergotropy exhibits non-trivial behavior during the environmental interaction. Namely, $\mathcal{R}_{iii}^{coh}$ increases over time, reaching a turning point at $\tau_{1/2}$ given by Eq. (31), when the diagonal state becomes passive. However, since $\tau_{\gamma}\approx\tau_{1/2}$ for the chosen parameters, the decaying part of the dynamics and the non-monotonicity are not observed, i.e., the reversal weak measurement is performed around the maximum value achieved by the coherent ergotropy during the thermalization. Also, under these conditions and in contrast to the incoherent part, the action of $\hat{M}_{m}$ and $\hat{W}_{\tilde{w}}$ respectively increase and decrease the coherent ergotropy. Despite that, at the end of the protocol $\mathcal{R}_{iv}^{coh}$ still remains above the red curve, resulting in a small finite gain of ¯$\overline{\mathcal{R}}^{coh}_{m\tilde{w}}(\tau_{\gamma})\approx 0.5\omega\times 10^{-2}$, equivalent to a $100\times\overline{\mathcal{R}}^{coh}_{m\tilde{w}}(\tau_{\gamma})/\mathcal{R}_{i}^{coh}\approx 5.66\%$ of extra charge.
Figure 8: Protocol performance in terms of the weak measurement strength $m$ and initial population $P_{0}$ for (a)-(c) $f=0.30$ and (d)-(f) $f=0.15$, higher and lower temperatures, respectively. These quantities were computed along the null energetic shift curve $\tilde{w}\in[0,1]$, considering $|Q_{0}|^{2}=|Q_{max}|^{2}$ for each $P_{0}$, and $\tau=\tau_{\gamma}$. The dashed blue curves indicate the points with null ergotropy gain, while the red lines highlight the points satisfying $\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$. (a) and (d) Total ergotropy gain; (b) and (e) Coherence contribution to the protocol, quantified by the coherent ergotropy gain; (c) and (f) Protocol’s success probability.
@@ III.3 Total ergotropy gain and success probability
Figure 9: Ergotropy gain $\times$ success probability in terms of $P_{0}$ and $m$, computed at $\tau_{\gamma}$ and $\tilde{w}\in[0,1]$ for (a)-(b) $f=0.30$ and (c)-(d) $f=0.15$. In (a) and (c) it was assumed $|Q_{0}|^{2}=|Q_{max}|^{2}$, while in (b) and (d) $|Q_{0}|^{2}=0$. The red dashed and continuous lines, and the hatched areas indicate the points where $\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$. The null energetic shift curve $\eta_{2}$ is highlighted.
@@
To better assess the overall performance of the protocol, Fig. 8 presents the relevant figures of merit for two different temperatures: higher temperature ($f=0.30$) in panels a–c and lower temperature ($f=0.15$) in panels d–f. Each row compares three key quantities as a function of the initial population $P_{0}$ and the weak measurement strength $m$. These quantities were computed along $\tilde{w}\in[0,1]$, dissipation time $\tau=\tau_{\gamma}$ and for the initial coherence it was assumed $|Q_{0}|^{2}=|Q_{max}|^{2}$ for each $P_{0}$. The blue dashed lines represent the points where the ergotropy gains are null, while the red continuous line highlights the protocol’s working points, where $\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$ is also satisfied. Figures 8(a) and (d) display the total ergotropy gain ¯$\overline{\mathcal{R}}_{m\tilde{w}}(\tau_{\gamma})$. Notice that, despite being positive for most of the surface $P_{0}\times m$, for the chosen set of parameters the operational points of the protocol are approximately constrained to a-c $P_{0}\in[0.6,0.85]$ for all $m$ and d-f $P_{0}\in[0.8,1]$ for high values of $m\in[0.95,1]$. Also, it is possible to observe that one obtains higher gains for stronger measurements, even in regions where the coherence contribution is negative. In this sense, Figures 8(b) and (e) show the coherent ergotropy gain ¯$\overline{\mathcal{R}}^{\text{coh}}_{m\tilde{w}}(\tau_{\gamma})$, revealing both positive and negative input in operational points, with stronger quantum coherence effects at lower temperatures (darker regions in panel (e)).
Figure 10: Ergotropy comparison. Ergotropy changes with (dashed blue) and without (continuous red) the execution of the protocol for $P_{0}=0.9$, $|Q_{0}|^{2}\approx 7.67\times 10^{-2}$, dissipation time $\tau=\tau_{\gamma}$, and strengths $m=0.4$ and $w=\tilde{w}\approx 0.2$. The blue arrows highlight the ergotropy jumps due to the measurements. The probability of success is $\Pi_{m\tilde{w}}(\tau_{\gamma})\approx 0.57$.
Finally, Figures 8(c) and (f) present the success probability $\Pi_{m\tilde{w}}(\tau_{\gamma})$, showing a very similar behavior for both the temperatures considered. As a probabilistic protocol, its performance should be evaluated alongside its probability of success. In this sense, note that ¯$\overline{\mathcal{R}}_{m\tilde{w}}(\tau_{\gamma})$ and $\Pi_{m\tilde{w}}(\tau_{\gamma})$ are inversely related in general, i.e., higher gains are obtained with lower probabilities of success and vice versa. Indeed, as described by Eq. (12), stronger measurements are less probable. On the other hand, as shown in panel (e), the coherent ergotropy gain may present a positive contribution in the higher probabilities region. It should be stressed that the observed probability landscape is a generic feature of the protocol, due to its independence of the initial coherence $Q_{0}$, i.e., the fact that the probability is maximized in the direction of lower $m$ and $P_{0}$ is observed regardless of the amount of coherence initially assumed. Overall, this figure provides a more detailed understanding of the protocol’s behavior, emphasizing that temperature influences both the total and coherent ergotropy gains, as well as the likelihood of success. Indeed, it is possible to observe that -for the chosen parameters- the operational points are constrained in the high-gain-low-probability region for lower temperatures (lower $f$), while for higher temperatures the number of operational points grows, suggesting that thermal fluctuations help balance the net ergotropy balance (see Appendix B).
Additionally, to better understand the relationship between gain and probability, Fig. 9 illustrates the ergotropy gain and the probability of success behavior assuming all values of $P_{0}$ and $m$ for $f=0.30$ in panels a-b and $f=0.15$ in panels c-d. Each row considers different values of initial coherence, with $|Q_{0}|^{2}=|Q_{max}|^{2}$ and $|Q_{0}|^{2}=0$ for the first and second row, respectively. The red lines and the hatched areas indicate the points where $\mathcal{W}_{m\tilde{w}}(\tau_{\gamma})=0$ is also satisfied. The red dots in (a) denote the working points with different values of $m$, while the red dashed line in (b) and (d) indicates the $\eta_{2}$ curve given by Eq. (25). The inverse relationship between ¯$\overline{\mathcal{R}}_{m\tilde{w}}(\tau_{\gamma})$ and $\Pi_{m\tilde{w}}(\tau_{\gamma})$ is evident and is due to the unlikely nature of strong measurements and the decay of the ergotropy during dissipation. When stronger measurements are applied (higher $m$ values), they create a more pronounced modification of the battery state, decreasing its ergotropy and leading to slower dissipation dynamics compared to the unmeasured case. Indeed, as discussed in Appendix A, the discharging rate for certain partially charged states is lower than that for charged ones. This enhanced state manipulation results in greater ergotropy gains by effectively modifying how the battery exchanges energy with its environment. However, these same stronger measurements substantially reduce the probability of obtaining the desired measurement outcomes, as they more definitively project the quantum state along specific measurement directions. This creates an inherent tradeoff, clearly illustrated in Fig. 9, where measurements with high strength values, e.g. $m=0.9$ produce substantial ergotropy gains but with low success probabilities. In contrast, weaker measurements ($m=0.1$) yield more modest gains with much higher reliability. This relationship suggests that practical implementations must balance effectiveness with reliability based on specific application requirements.